OpenStudy (anonymous):

Please Help...Statistics! 52% of men consider themselves professional baseball fans. You randomly select 10 men & ask each if he considers himself a professional baseball fan. Find the probability that the number who consider themselves baseball fan is a) exactly 8 b) at least 8 c) less than 8 If convenient use technology to find the probabilities.

OpenStudy (kropot72):

a) The binomial distribution applies. The probability of exactly 8 fans out of 10 is: \[P(8:10)=\left(\begin{matrix}10 \\ 8\end{matrix}\right)\times (0.52)^{8}\times (0.48)^{2}\] What is your answer?

OpenStudy (kropot72):

Do you need help with the calculation?

OpenStudy (kropot72):

@Solunus Are you there?

OpenStudy (anonymous):

yeah, that's it. way better at this stuff than I.

OpenStudy (kropot72):

\[\left(\begin{matrix}10 \\ 8\end{matrix}\right)=\frac{10\times 9}{2}=45\] So the calculation is as follows: \[45\times (0.52)^{8}\times (0.48)^{2}=?\]